# GOLD: Graph Out-of-Distribution Detection via Implicit Adversarial Latent Generation

### v1

@@ 2 Preliminary
To detect the testing nodes coming from a different distribution from the training data, where $P_{train}(\mathbf{X,A})\neq P_{test}(\mathbf{X,A})$ and the conditional distribution $P_{train}(\mathbf{y}|\mathbf{X,A})\neq P_{test}(\mathbf{y}|\mathbf{X,A})$, the task is to require an OOD detector $F$ to output a binary prediction for the testing nodes. $F$ is usually built upon the output from the classifier GNN with $F(\mathbf{X},\mathbf{A};\text{GNN})=0$ for data from in-distribution and $F(\mathbf{X},\mathbf{A};\text{GNN})=1$ for data from out-of-distribution (Wu et al., 2023b; Liu et al., 2020; Yang et al., 2023a).
@@ 3.2 OOD Detector for ID and Pseuro-OOD Separation
$\max_{\text{GNN, MLP}}\mathcal{L}_{\text{Unc}},\text{ where }\mathcal{L}_{\text{Unc}}=\mathbb{E}_{i\sim P_{\text{ID}}}\log[\phi(\text{MLP}(e_{i})_{{[0]}})]+\mathbb{E}_{j\sim P_{\text{p-OOD}}}\log[\phi(\text{MLP}(e_{j})_{{[1]}})].$ (10)
@@ Appendix A Appendix
$\begin{split}-\frac{\partial\mathcal{L}_{\text{Unc}}}{\partial\theta}&=-\mathbb{E}_{i\sim P_{\text{ID}}}\frac{\partial\log[\phi(l_{\theta}(e_{i})_{[0]})]}{\partial\theta}-\mathbb{E}_{j\sim P_{\text{p-OOD}}}\frac{\partial\log[\phi(l_{\theta}(e_{j})_{[1]})]}{\partial\theta}\\ &=-\mathbb{E}_{i\sim P_{\text{ID}}}\frac{\partial\log[\frac{e^{l_{\theta}(e_{i})_{[0]}}}{e^{l_{\theta}(e_{i})_{[0]}}+e^{l_{\theta}(e_{i})_{[1]}}}]}{\partial\theta}-\mathbb{E}_{j\sim P_{\text{p-OOD}}}\frac{\partial\log[\frac{e^{l_{\theta}(e_{j})_{[1]}}}{e^{l_{\theta}(e_{j})_{[0]}}+e^{l_{\theta}(e_{j})_{[1]}}}]}{\partial\theta}\\ &=-\mathbb{E}_{i\sim P_{\text{ID}}}\frac{\partial[l_{\theta}(e_{i})_{[0]}-\log(e^{l_{\theta}(e_{i})_{[0]}}+e^{l_{\theta}(e_{i})_{[1]}})]}{\partial\theta}\\ &-\mathbb{E}_{j\sim P_{\text{p-OOD}}}\frac{[\partial l_{\theta}(e_{j})_{[1]}-\log(e^{l_{\theta}(e_{j})_{[0]}}+e^{l_{\theta}(e_{j})_{[1]}})]}{\partial\theta}\\ &=\mathbb{E}_{i\sim P_{\text{ID}}}\left[-\frac{\partial l_{\theta}(e_{i})_{[0]}}{\partial\theta}+\frac{e^{\l_{\theta}(e_{i})_{[0]}}\frac{\partial\l_{\theta}(e_{i})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{i})_{[1]}}\frac{\partial\l_{\theta}(e_{i})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{i})_{[0]}}+e^{\l_{\theta}(e_{i})_{[1]}}}\right]\\ &+\mathbb{E}_{j\sim P_{\text{p-OOD}}}\left[-\frac{\partial l_{\theta}(e_{j})_{[1]}}{\partial\theta}+\frac{e^{\l_{\theta}(e_{j})_{[0]}}\frac{\partial\l_{\theta}(e_{j})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{j})_{[1]}}\frac{\partial\l_{\theta}(e_{j})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{j})_{[0]}}+e^{\l_{\theta}(e_{j})_{[1]}}}\right]\end{split}$ (16)
@@
$\begin{split}-(\frac{\partial\mathcal{L}_{\text{Unc}}}{\partial\theta}+\frac{\partial\mathcal{L}_{\text{DReg}_{S}}}{\partial\theta})&=\mathbb{E}_{i\sim P_{\text{ID}}}\left[-\frac{\partial l_{\theta}(e_{i})_{[0]}}{\partial\theta}+\frac{e^{\l_{\theta}(e_{i})_{[0]}}\frac{\partial\l_{\theta}(e_{i})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{i})_{[1]}}\frac{\partial\l_{\theta}(e_{i})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{i})_{[0]}}+e^{\l_{\theta}(e_{i})_{[1]}}}\right]\\ &+\mathbb{E}_{j\sim P_{\text{p-OOD}}}\left[-\frac{\partial l_{\theta}(e_{j})_{[1]}}{\partial\theta}+\frac{e^{\l_{\theta}(e_{j})_{[0]}}\frac{\partial\l_{\theta}(e_{j})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{j})_{[1]}}\frac{\partial\l_{\theta}(e_{j})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{j})_{[0]}}+e^{\l_{\theta}(e_{j})_{[1]}}}\right]\\ &-\mathbb{E}_{i\sim P_{\text{ID}}}\frac{e^{\l_{\theta}(e_{i})_{[0]}}\frac{\partial\l_{\theta}(e_{i})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{i})_{[1]}}\frac{\partial\l_{\theta}(e_{i})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{i})_{[0]}}+e^{\l_{\theta}(e_{i})_{[1]}}}\\ &+\mathbb{E}_{j\sim P_{\text{p-OOD}}}\frac{e^{\l_{\theta}(e_{j})_{[0]}}\frac{\partial\l_{\theta}(e_{j})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{j})_{[1]}}\frac{\partial\l_{\theta}(e_{j})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{j})_{[0]}}+e^{\l_{\theta}(e_{j})_{[1]}}}\\ &\text{Define the energy w.r.t. label y as $E(e,y)=-l_{\theta}(e)_{[y]}$}\\ &=\mathbb{E}_{i\sim P_{\text{ID}}}\frac{\partial E(e_{i},0)}{\partial\theta}+\mathbb{E}_{j\sim P_{\text{p-OOD}}}\frac{\partial E(e_{j},1)}{\partial\theta}\\ &-2\left(\phi(\l_{\theta}(e_{j})_{[0]})\frac{\partial E(e_{j},0)}{\partial\theta}+\phi(\l_{\theta}(e_{j})_{[1]})\frac{\partial E(e_{j},1)}{\partial\theta}\right)\\ &=\mathbb{E}_{i\sim P_{\text{ID}}}\frac{\partial E(e_{i},0)}{\partial\theta}+\mathbb{E}_{j\sim P_{\text{p-OOD}}}(1-2\phi(\l_{\theta}(e_{j})_{[1]}))\frac{\partial E(e_{j},1)}{\partial\theta}\\ &-2\phi(\l_{\theta}(e_{j})_{[0]})\frac{\partial E(e_{j},0)}{\partial\theta}\end{split}$ (18)
From the above equation, the training procedure that overall minimises the first order gradient of the negative sum of $\mathcal{L}_{\text{Unc}}$ and the surrogate function of $\mathcal{L}_{\text{DReg}}$ will decrease the energy score $E(e_{i},0;l_{\theta})$ for in-distribution data, and increase the energy score $E(e_{j},1;l_{\theta})$ and $E(e_{j},0;l_{\theta})$ for pseudo-OOD data, given $\phi(\l_{\theta}(e_{j})_{[1]})>0.5$ as the detector continues to improve detection performance. ∎

### v2

@@ 2 Preliminary
To detect the testing nodes coming from a different distribution from the training data, where $P_{train}(\mathbf{X,A})\neq P_{test}(\mathbf{X,A})$ and the conditional distribution $P_{train}(\mathbf{y}|\mathbf{X,A})\neq P_{test}(\mathbf{y}|\mathbf{X,A})$, the task is to require an OOD detector $F$ to output a binary prediction for the testing nodes. $F$ is usually built upon the output from the classifier GNN with $F(\mathbf{x},\mathbf{A};\text{GNN})=1$ for data from in-distribution and $F(\mathbf{x},\mathbf{A};\text{GNN})=0$ for data from out-of-distribution (Wu et al., 2023b; Liu et al., 2020; Yang et al., 2023a).
@@ 3.2 OOD Detector for ID and Pseuro-OOD Separation
$\max_{\text{GNN, MLP}}\mathcal{L}_{\text{Unc}},\text{ where }\mathcal{L}_{\text{Unc}}=\mathbb{E}_{i\sim P_{\text{ID}}}\log[\phi(\text{MLP}(e_{i}))_{{[0]}}]+\mathbb{E}_{j\sim P_{\text{p-OOD}}}\log[\phi(\text{MLP}(e_{j}))_{{[1]}}].$ (10)
@@ Appendix A Appendix
$\begin{split}-\frac{\partial\mathcal{L}_{\text{Unc}}}{\partial\theta}&=-\mathbb{E}_{i\sim P_{\text{ID}}}\frac{\partial\log[\phi(l_{\theta}(e_{i}))_{[0]}]}{\partial\theta}-\mathbb{E}_{j\sim P_{\text{p-OOD}}}\frac{\partial\log[\phi(l_{\theta}(e_{j}))_{[1]}]}{\partial\theta}\\ &=-\mathbb{E}_{i\sim P_{\text{ID}}}\frac{\partial\log[\frac{e^{l_{\theta}(e_{i})_{[0]}}}{e^{l_{\theta}(e_{i})_{[0]}}+e^{l_{\theta}(e_{i})_{[1]}}}]}{\partial\theta}-\mathbb{E}_{j\sim P_{\text{p-OOD}}}\frac{\partial\log[\frac{e^{l_{\theta}(e_{j})_{[1]}}}{e^{l_{\theta}(e_{j})_{[0]}}+e^{l_{\theta}(e_{j})_{[1]}}}]}{\partial\theta}\\ &=-\mathbb{E}_{i\sim P_{\text{ID}}}\frac{\partial[l_{\theta}(e_{i})_{[0]}-\log(e^{l_{\theta}(e_{i})_{[0]}}+e^{l_{\theta}(e_{i})_{[1]}})]}{\partial\theta}\\ &-\mathbb{E}_{j\sim P_{\text{p-OOD}}}\frac{[\partial l_{\theta}(e_{j})_{[1]}-\log(e^{l_{\theta}(e_{j})_{[0]}}+e^{l_{\theta}(e_{j})_{[1]}})]}{\partial\theta}\\ &=\mathbb{E}_{i\sim P_{\text{ID}}}\left[-\frac{\partial l_{\theta}(e_{i})_{[0]}}{\partial\theta}+\frac{e^{\l_{\theta}(e_{i})_{[0]}}\frac{\partial\l_{\theta}(e_{i})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{i})_{[1]}}\frac{\partial\l_{\theta}(e_{i})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{i})_{[0]}}+e^{\l_{\theta}(e_{i})_{[1]}}}\right]\\ &+\mathbb{E}_{j\sim P_{\text{p-OOD}}}\left[-\frac{\partial l_{\theta}(e_{j})_{[1]}}{\partial\theta}+\frac{e^{\l_{\theta}(e_{j})_{[0]}}\frac{\partial\l_{\theta}(e_{j})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{j})_{[1]}}\frac{\partial\l_{\theta}(e_{j})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{j})_{[0]}}+e^{\l_{\theta}(e_{j})_{[1]}}}\right]\end{split}$ (16)
@@
$\begin{split}-(\frac{\partial\mathcal{L}_{\text{Unc}}}{\partial\theta}+\frac{\partial\mathcal{L}_{\text{DReg}_{S}}}{\partial\theta})&=\mathbb{E}_{i\sim P_{\text{ID}}}\left[-\frac{\partial l_{\theta}(e_{i})_{[0]}}{\partial\theta}+\frac{e^{\l_{\theta}(e_{i})_{[0]}}\frac{\partial\l_{\theta}(e_{i})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{i})_{[1]}}\frac{\partial\l_{\theta}(e_{i})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{i})_{[0]}}+e^{\l_{\theta}(e_{i})_{[1]}}}\right]\\ &+\mathbb{E}_{j\sim P_{\text{p-OOD}}}\left[-\frac{\partial l_{\theta}(e_{j})_{[1]}}{\partial\theta}+\frac{e^{\l_{\theta}(e_{j})_{[0]}}\frac{\partial\l_{\theta}(e_{j})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{j})_{[1]}}\frac{\partial\l_{\theta}(e_{j})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{j})_{[0]}}+e^{\l_{\theta}(e_{j})_{[1]}}}\right]\\ &-\mathbb{E}_{i\sim P_{\text{ID}}}\frac{e^{\l_{\theta}(e_{i})_{[0]}}\frac{\partial\l_{\theta}(e_{i})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{i})_{[1]}}\frac{\partial\l_{\theta}(e_{i})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{i})_{[0]}}+e^{\l_{\theta}(e_{i})_{[1]}}}\\ &+\mathbb{E}_{j\sim P_{\text{p-OOD}}}\frac{e^{\l_{\theta}(e_{j})_{[0]}}\frac{\partial\l_{\theta}(e_{j})_{[0]}}{\partial\theta}+e^{\l_{\theta}(e_{j})_{[1]}}\frac{\partial\l_{\theta}(e_{j})_{[1]}}{\partial\theta}}{e^{\l_{\theta}(e_{j})_{[0]}}+e^{\l_{\theta}(e_{j})_{[1]}}}\\ &\text{Define the energy w.r.t. label y as $E(e,y)=-l_{\theta}(e)_{[y]}$}\\ &=\mathbb{E}_{i\sim P_{\text{ID}}}\frac{\partial E(e_{i},0)}{\partial\theta}+\mathbb{E}_{j\sim P_{\text{p-OOD}}}\frac{\partial E(e_{j},1)}{\partial\theta}\\ &-2\left(\phi(\l_{\theta}(e_{j}))_{[0]}\frac{\partial E(e_{j},0)}{\partial\theta}+\phi(\l_{\theta}(e_{j}))_{[1]}\frac{\partial E(e_{j},1)}{\partial\theta}\right)\\ &=\mathbb{E}_{i\sim P_{\text{ID}}}\frac{\partial E(e_{i},0)}{\partial\theta}+\mathbb{E}_{j\sim P_{\text{p-OOD}}}(1-2\phi(\l_{\theta}(e_{j}))_{[1]})\frac{\partial E(e_{j},1)}{\partial\theta}\\ &-2\phi(\l_{\theta}(e_{j}))_{[0]}\frac{\partial E(e_{j},0)}{\partial\theta}\end{split}$ (18)
From the above equation, the training procedure that overall minimises the first order gradient of the negative sum of $\mathcal{L}_{\text{Unc}}$ and the surrogate function of $\mathcal{L}_{\text{DReg}}$ will decrease the energy score $E(e_{i},0;l_{\theta})$ for in-distribution data, and increase the energy score $E(e_{j},1;l_{\theta})$ and $E(e_{j},0;l_{\theta})$ for pseudo-OOD data, given $\phi(\l_{\theta}(e_{j}))_{[1]}>0.5$ as the detector continues to improve detection performance. ∎